# Experimental Tests of Invariant Set Theory

### v2

@@ V Gravitational Decoherence
Figure 4: The experiment described by Bose et al [35] and Marletto and Vedral [36], for testing gravity’s ability to entangle two masses. Two masses, $m_{i}$ for $i\in\{1,2\}$ are separated from each other by distance $d$. Both are initially in state $\left|C\right\rangle_{i}$, with embedded spin $(\left|\uparrow\right\rangle+\left|\downarrow\right\rangle)/\sqrt{2}$. They are then both admitted into Stern-Gerlach devices, which put them both into the spin-dependent superposition $(\left|L,\uparrow\right\rangle_{i}+\left|R,\downarrow\right\rangle_{i})/\sqrt{2}$, where $\left|L\right\rangle_{i}$ and $\left|R\right\rangle_{i}$ are separated from each other by distance $\Delta x_{i}$. They are left in these superpositions for time $\tau$. During this time, if gravity is quantum-coherent, evolution under mutual gravitational attraction $h_{00}$ would entangle the two particles, adding relevant phases to both. After time $\tau$, an inverse Stern-Gerlach device is applied to return each mass to their initial state (potentially modulo the phases applied by $h_{00}$). By applying this process, and measuring spin correlations between the two particles after each run, we can detect if relative phases have been applied to each, and so if gravity is coherent. For IST to hold, gravity must be decoherent, and so cannot entangle two masses. This means IST predicts no alteration of phases will be detected.
Palmer describes IST as not so much a quantum theory of gravity (like String Theory and Loop Quantum Gravity), but a gravitational theory of the quantum [7]. Aside from its determinate nature, nowhere is this more true than in how IST models regimes where gravitational and quantum effects are both present. The paper Invariant Set Theory describes the theory as positing no gravitons and so no supersymmetry (spin-2 gravitons typically being seen as hinting at supersymmetry) [7]. Instead, the paper holds that gravity is inherently decoherent, turning gravitationally-affected superpositions into maximally mixed states. This paper also claims that effects typically considered signs of either dark matter or dark energy could instead be in some way due to various manifestations of the “smearing” of energy-momentum on space-times neighbouring our universe $\mathcal{M}_{U}$ on the invariant fractal set $I_{U}$ influencing curvature of $\mathcal{M}_{U}$. It claims this smearing avoids precise singularities in $\mathcal{M}_{U}$: avoiding singularities being a key goal of many previous attempts to quantise General Relativity.
Palmer suggests an alteration of the Einstein Field Equation (EFE) based on the presence and effects of possible universes $\mathcal{M}^{\prime}_{U}$ on our universe $\mathcal{M}_{U}$, leading to the EFE instead being
@@
This gravitational decoherence could be tested by experiments that involve putting heavy objects in spatial superpositions. This would involve allowing them to gravitationally interact, then returning the spatial superposition components back to a single position, then seeing if there are any signs of entanglement between the objects from the resulting interference pattern (see Fig.4) [35, 36].
@@
While such a test may sound challenging to implement, recent experimental work indicates it should be possible within the next few years [37, 38]. This test would also have the benefit of either ruling for or ruling out other theories which hold gravity to be decoherent (e.g. Penrose’s gravitational collapse interpretation [39].
VI Conclusions
@@
We have identified points of difference between Invariant Set Theory and standard quantum theory. While these are not fatal to IST, they provide potential avenues to experimentally test the theory, to see whether its deterministic, fractal-attractor-based structure is compatible with observed reality. We described differences between the empirical predictions of IST and standard quantum mechanics. We then proposed experiments, based on current or near-future quantum technologies (e.g. noisy intermediate-scale quantum devices), which would test which set of predictions most closely matches reality. This serves as an example of how these near-term quantum technologies allow us to probe into the foundational mysteries of quantum mechanics.
@@ Acknowledgements.
JRH and JR acknowledge support from the University of York’s EPSRC DTP grant EP/R513386/1, and the Quantum Communications Hub funded by the EPSRC grants EP/M013472/1 and EP/T001011/1. TP acknowledges support from a Royal Society Research Professorship. JRH acknowledges support from Hiroshima University’s Phoenix Postdoctoral Fellowship for Research programme.

### v3

@@ V Gravitational Decoherence
Figure 4: The experiment described by Bose et al [31] and Marletto and Vedral [32], to test whether gravity can entangle two masses. Two masses, $m_{i}$ for $i\in\{1,2\}$ are separated from each other by distance $d$. Both are initially in state $\left|C\right\rangle_{i}$, with embedded spin $(\left|\uparrow\right\rangle+\left|\downarrow\right\rangle)/\sqrt{2}$. They are then both admitted into Stern-Gerlach devices, which put them both into the spin-dependent superposition $(\left|L,\uparrow\right\rangle_{i}+\left|R,\downarrow\right\rangle_{i})/\sqrt{2}$, where $\left|L\right\rangle_{i}$ and $\left|R\right\rangle_{i}$ are separated from each other by distance $\Delta x_{i}$. They are left in these superpositions for time $\tau$. During this time, if gravity is quantum-coherent, evolution under mutual gravitational attraction $h_{00}$ would entangle the two particles, adding relevant phases to both. After time $\tau$, an inverse Stern-Gerlach device is applied to return each mass to their initial state (potentially modulo the phases applied by $h_{00}$). By applying this process, and measuring spin correlations between the two particles after each run, we can detect if relative phases have been applied to each, and so if gravity is coherent. For IST to hold, gravity must be decoherent, and so cannot entangle two masses. This means IST predicts no alteration of phases will be detected.
Palmer describes IST as not so much a quantum theory of gravity (like String Theory and Loop Quantum Gravity), but a gravitational theory of the quantum [3]. Aside from its determinate nature, nowhere is this more true than in how IST models regimes where gravitational and quantum effects are both present. The paper Invariant Set Theory describes the theory as positing no gravitons and so no supersymmetry [3] (spin-2 gravitons typically being seen as hinting at supersymmetry [33]). Instead, the paper holds that gravity is inherently decoherent, turning gravitationally-affected superpositions into maximally mixed states. This paper also claims that effects typically considered signs of either dark matter or dark energy could instead be in some way due to various manifestations of the “smearing” of energy-momentum on space-times neighbouring our universe $\mathcal{M}_{U}$ on the invariant fractal set $I_{U}$ influencing curvature of $\mathcal{M}_{U}$. It claims this smearing avoids precise singularities in $\mathcal{M}_{U}$: avoiding singularities being a key goal of many previous attempts to quantise General Relativity.
Palmer suggests an alteration of the Einstein Field Equation (EFE) [34] based on the presence and effects of possible universes $\mathcal{M}^{\prime}_{U}$ on our universe $\mathcal{M}_{U}$, leading to the EFE instead being
@@
This gravitational decoherence could be tested by experiments that involve putting heavy objects in spatial superpositions. This would involve allowing them to gravitationally interact, then returning the spatial superposition components back to a single position, then seeing if there are any signs of entanglement between the objects from the resulting interference pattern (see Fig. 4) [31, 32, 35, 36].
@@
While such a test may sound challenging to implement, recent experimental work indicates it should be possible within the next few years [37, 38, 39, 40, 41]. This test would also have the benefit of either ruling for or ruling out other theories which hold gravity (or some other measure of macroscopicity) to be decoherent: e.g. the Penrose-Diósi gravitational collapse model [42, 43, 44], or other spontaneous collapse models, such as the continuous spontaneous localisation (CSL) model [45, 46] or the the Ghirardi-Rimini-Weber (GRW) model [47]. While these models may seem different to IST, they have one key thing in common with it—none of them hold quantum mechanics to be complete.
VI Discussion
Unfortunately, all of the tests we discuss above suffer from one key limitation in their ability to discriminate the experimental predictions of Invariant Set Theory from standard quantum mechanics: the fact that $p$ has not yet been bounded in the literature. Note though this is not unusual in the literature, similarly to the difficulties bounding the phenomenological parameters $\lambda$ and $r_{C}$ for the Ghirardi-Rimini-Weber [47] and Continuous Spontaneous Localisation [45] spontaneous collapse models, or the mass-density distribution $R_{0}$ for the Diósi-Penrose gravitational collapse model [42, 43]. Identifying the key role played by these free parameters in experimentally differentiating these models from standard quantum mechanics has led to a strong focus in bounding their allowed values from different directions [48, 49, 50, 51], a key aspect of the spontaneous collapse model research programme. We hope, by pointing out that the prospect of finding experimental tests for Invariant Set theory rests heavily on being able to bound $p$, this work will motivate interested researchers to identify similar bounds as has been done for the free parameters of these other models.
One question the more reader might have, after reading this paper, is why we have gone to the lengths of considering experimental tests for one specific model, arguably developed enough to only constitute a toy model (rather than a full theory describing and extending beyond quantum mechanics). Our answer is that the meaningfulness of physical models arguably comes from the empirical predictions they make, and specifically the differences in the empirical predictions they make from what is commonly accepted (here, standard quantum mechanics), rather than just from their mathematical structure. We can see this for standard quantum mechanics itself, where the meaningfulness of the completeness/incompleteness debate arguably didn’t come from the discussions of Bohr and Einstein [52, 53, 54], but from Bell’s hypothesis that there would be experimental differences between approaches where quantum mechanics was considered complete (i.e., standard quantum mechanics), and certain approaches where it was considered incomplete (specifically, local hidden variable models respecting statistical independence between measurement settings and hidden variables) [55, 56], and Clauser et al’s attempts to derive experimental proposals where such a difference was demonstrated [57, 58]. The fact that the burgeoning field of quantum technologies did not begin at the advent of quantum mechanics, but after standard quantum mechanics had been differentiated from these models is testament to this idea. Elsewhere, we can see this in the rise and fall of String Theory—where interesting mathematical structures attracted many young researchers, the lack of any empirically-testable predictions (and several popular books pointing this out [59, 60]) has left the field slowly wilting away into obscurity. Empirical validity or falsifiability is key for new hypotheses to be meaningful, to attract attention, and ultimately to be usable to develop useful new tools—without this, the hypothesis is arguably just abstract theory or metaphysics. Given (despite not yet being a complete theory) Invariant Set Theory is the archetype for supermeasured theories [11, 61] (those which nominally violate the assumption of statistical independence in Bell’s Theorem in a different way to how superdeterministic or retrocausal models are typically assumed to), looking at ways to make IST more physically meaningful is a key research priority for those of us interested in models violating statistical independence. We hope this work, for all its limitations, has succeeded in increasing the meaningfulness of this model.
@@ VI Discussion
To conclude, we have identified points of difference between Invariant Set Theory and standard quantum theory. While these are not fatal to IST, they provide potential avenues to experimentally test the theory, to see whether its deterministic, fractal-attractor-based structure is compatible with observed reality. We described differences between the empirical predictions of IST and standard quantum mechanics. We then proposed experiments, based on current or near-future quantum technologies (e.g. noisy intermediate-scale quantum devices), which would test which set of predictions most closely matches reality. This serves as an example of how these near-term quantum technologies allow us to probe into the foundational mysteries of quantum mechanics.
@@ Acknowledgements.
JRH acknowledges support from the University of York’s EPSRC DTP grant EP/R513386/1, and Hiroshima University’s Phoenix Postdoctoral Fellowship for Research programme. JRH and JR acknowledge support from the Quantum Communications Hub funded by the EPSRC grants EP/M013472/1 and EP/T001011/1. TP acknowledges support from a Royal Society Research Professorship.